# Beyond Trend and Periodicity: Guiding Time Series Forecasting with Textual Cues

### v2

@@ 1 Introduction
Time series forecasting (TSF) is crucial in many fields, thereby attracting significant attention from both academia and industry. Despite extensive research dedicated to this task, recent studies have shown that simple linear models (Zeng et al., 2023; Xu et al., 2023; Toner & Darlow, 2024) that barely extract trend and periodicity information from time series data frequently achieve performance close to that of state-of-the-art complex models (Nie et al., 2023; Liu et al., 2023; Jin et al., 2023), indicating that current approaches have reached a saturation point, possibly overfitting historical data.
To achieve higher forecasting accuracy, it is essential to provide models with additional, critical information—such as holiday information, consumer sentiment, and climate changes (depending on the specific forecasting task)—that are not present in the time series data itself. Without this additional information, models cannot extract the necessary causal relationships, leading to oversimplification or overfitting. Moreover, the lack of insight about channels and underlying systems prevents the modeling of inter-channel relationships, often resulting in unreliable forecasts.
As such valuable information is usually hard to quantify as auxiliary time series data, we consider them as textual cues and introduce a novel task for TSF: Text-Guided Time Series Forecasting (TGTSF). TGTSF leverages two main components: channel descriptions and dynamic news messages. Channel descriptions provide static knowledge about the underlying systems, enabling the model to differentiate between channels and better understand inter-channel correlations. News messages offer dynamic and external insights, helping the model adjust to shifts in data distribution caused by external events. By modeling the joint distribution of future values conditioned on these textual cues, TGTSF has the potential to enhance forecasting accuracy and reliability. Furthermore, TGTSF enables scenario-based forecasting, which is particularly valuable for business decision-making. For instance, a company can forecast sales based on the assumption of a successful marketing campaign, incorporating relevant news and channel-specific descriptions to refine predictions.
We then introduce TGForecaster, a simple Transformer-based baseline multimodal model designed to realize the potential of the proposed TGTSF framework. Utilizing a cross-attention mechanism, TGForecaster fuses information from textual cues and time series data to guide the forecasting process. To rigorously evaluate and demonstrate the efficacy of our model, we carefully craft four benchmark datasets categorized into synthetic, captioned existing, and real-world datasets. Each dataset is uniquely designed to validate the task and test the model’s capability from different perspectives. Specifically, the synthetic dataset features simple yet variable time series patterns, such as sinusoidal waves with random frequency changes, to test the model’s ability to utilize textual cues in forecasting and mitigating distribution shifts. Two existing datasets are captioned with different levels of textual information, used to challenge the model in complex, non-periodic environments by integrating detailed external data. Lastly, our real-world dataset, collected from actual sales scenarios, showcases the practical application value of TGTSF.
Experimental results show that TGForecaster consistently demonstrates state-of-the-art performance across our datasets, affirming the soundness of the TGTSF task definition. Ablation studies reveal that without textual assistance, TGForecaster’s performance reverts to that of PatchTST (Nie et al., 2023), its time series encoder backbone, underscoring that the performance enhancement is driven by the additional information provided by textual data, not merely by a more sophisticated architecture.
Our contributions are summarized as follows:
  • We identify the roadblock of TSF as information insufficiency and propose TGTSF as a new forecasting approach that integrates textual data to enrich the models with external causal information and system knowledge.
  • We establish the first TGTSF benchmark containing four uniquely designed datasets.
  • We design a simple baseline model for TGTSF, TGForecaster. TGForecaster achieves state-of-the-art performance by utilizing textual information and effectively validates our proposed TGTSF task.
2 Insights and Motivation
2.1 Existing TSF Solutions Suffer from Information Insufficiency
Time series forecasting typically involves predicting future segments based on historical data, specifically forecasting subsequent time series segments from prior ones. However, the inherent sparsity of information within time series data makes achieving accurate forecasts challenging, or even impossible at times. Standard decomposition of time series data identifies three main components: trend, periodicity, and noise (Box et al., 2015). The noise component is intrinsically unpredictable. Trends, which are slowly changing patterns, are often impacted by external events that influence the underlying system, making it less predictable without external information. As a result, the periodicity component is usually the main source of reliable prediction.
Recent advancements in linear models have validated these challenges. For instance, the DLinear (Zeng et al., 2023) model utilizes a simple linear layer to effectively capture basic periodic patterns, outperforming most complex Transformer-based models. Similarly, FITS (Xu et al., 2023), designed to simply extract periodicity, achieves comparable or superior results to state-of-the-art (SOTA) methods using fewer than $10k$ parameters. Further research demonstrates that employing a closed-form forecasting matrix (Toner & Darlow, 2024), calculated from training data, can produce SOTA outcomes. These findings indicate that the minimal information in time series data allows models to recognize patterns with few or no learnable parameters, suggesting that training larger models with insufficient information can lead to severe overfitting.
The issue of information insufficiency extends beyond the data itself to the absence of external information. External factors like holidays, consumer sentiment, or climate changes can significantly alter patterns in time series data, affecting variables such as electricity usage, sales trends, or regional weather conditions. However, these factors are not contained within the time series data, hindering models from incorporating causal relationships. This deficiency can compromise model training, causing models to either learn overly simplified average shortcut or overfit on the training data, as shown in Fig. 1. For example, if a dataset frequently presents days without rain, the model may consistently predict no rain for future days. However, rainfall predictions are largely influenced by external factors like climate change, not included in the time series, leading models to opt for shortcuts that minimize overall loss by predicting the most common outcome.
Moreover, the lack of knowledge about the underlying systems exacerbates the problem of information insufficiency, especially in datasets with multiple time series channels. Without a deeper understanding of these systems or the characteristics of each channel, it is challenging to leverage such information for improved modeling performance. The indistinctiveness among channels complicates modeling efforts, making it difficult to accurately capture and model correlations.
These challenges underscore the need to integrate external information to address the fundamental limitations of time series forecasting methods.
2.2 RIN & Weight Sharing are Compromises under Information Insufficiency
Researchers have developed various methods to tackle information insufficiency in time series forecasting. One such method is Reversible Instance Normalization (RIN) (Kim et al., 2021), which normalizes each instance to align mean and variance across the dataset, as shown in Fig. 1. RIN effectively reduces distribution shifts caused by trends shifting or external events, forcing the model to learn a unified distribution. While it standardizes data inputs, RIN also strips away essential information such as trend intensity and amplitude variations, which diminishes the model’s capacity to detect relative biases or amplitude shifts beyond the training scope. However, RIN merely hides the issue of information insufficiency by neutralizing the effects of external events.
Figure 1: Left Panel: The traditional TSF model, when faced with distribution shifts due to external events, tends to learn an averaged shortcut for a smaller global loss, as shown in the leftmost figure. Conversely, the TGTSF task utilizes text inputs to eliminate the uncertainty and effectively generalize to these shifts, as depicted in the second figure. Right Panel: The rightmost figure illustrates the effect of text guidance on adapting to different channel distributions. Weight Sharing (WS) method helps the model capture only common knowledge among channels, while RIN addresses distribution shifts by normalizing instances, thus learning a larger portion of the distribution. TGTSF, however, leverages text to directly learn the entire distribution, enabling the model to generalize to a specific channel distribution with corresponding text as the condition.
Another strategy is weight sharing (Nie et al., 2023; Zeng et al., 2023; Xu et al., 2023). This technique treats all input channels as equivalent, leveraging common patterns across all channels to alleviate the problem of limited training samples. In other words, this approach enhances performance by learning from the shared periodicities across all the channels, typical within a single dataset. However, since the model is trained to minimize loss across all channel distributions with a single set of weights, it prioritizes common patterns, failing to capture unique distributions for each channel, as illustrated in Fig. 1, which may lead to averaged results. These average results further reveal that the model lacks sufficient information to distinguish between different channels effectively.
Despite these advancements, both methods provide only temporary solutions to the pervasive issue of information insufficiency. They do not resolve the core problem and still suffer from the limitations imposed by insufficient information.
2.3 Motivation for The Proposed TGTSF Framework
External factors significantly impact time series behavior such as trending and periodicity. A line of work apply and investigate how to embed the external information as a new channel, such as Autoformer (Wu et al., 2021), FedFormer (Zhou et al., 2022a) add auxilary channels to indicate the day-of-week or even the exact date. Temporal Fusion Transformer (TFT) (Lim et al., 2021) and TimeXer (Wang et al., 2024), however, focus on how to select the given variables and how to embed and fuse them with other channels. However, these methods only bring limited performance boost. Basically, they are still trying to squeeze out the information in the original time series dataset which contains limited information itself.
However, other forms of information such as news articles, social media updates, global climate change report are crucial for accurately predicting time series pattern in domains like sales and local weather report. These information are often in forms of text and hard to be quantified as numerical data alone. Moreover, the real-world application of forecasting frequently requires conditional predictions based on specific scenarios. For example, businesses may need to assess the impact of a marketing campaign on future sales, necessitating a deep understanding of external textual factors to forecast outcomes effectively.
To address these challenges, it is essential to include text as an additional input modality. This approach not only allows models to react to historical data but also to proactively adapt to potential future events and trends. It enhances forecasting by providing insights into the underlying dynamics and causal relationships often missing from time series data alone. Guiding time series forecasting with textual data also helps clarify inter-channel relationships, reflecting systemic correlations or domain-specific knowledge that are crucial for accurate predictions. This textual context enhances models’ ability to interpret complex data interactions, vital in environments where data channels are interconnected.
In summary, guiding time series forecasting with textual data addresses the critical issue of information insufficiency. This approach transcends the limitations of traditional TSF methodologies, enhancing model accuracy and effectiveness. By incorporating textual insights, forecasting becomes more robust and practically valuable for real-world applications.
3 TGTSF Task Formulation
We introduce a novel task within the domain of time series analysis, termed Text-Guided Time Series Forecasting (TGTSF), designed to address the prevalent issue of information insufficiency in time series forecasting.
As illustrated in Fig. 1, text guidance in TGTSF operates on two aspects. Firstly, channel descriptions serve as identifiers for each channel, aiding the model in distinguishing between them while learning shared features. These descriptions can also incorporate static knowledge about the underlying system, enhancing the model’s ability to recognize inter-channel correlations. Secondly, news messages provide dynamic and external insights into known (in training) or hypothesized (in inference) future events, which assist the model in adapting to event-driven distribution shifts.  [^1] In creating a TGTSF dataset, it is advisable to avoid directly generating the description out of the forecasting horizon time series pattern as news messages, as this could lead to information leakage. News messages should instead contain relevant, known information from other sources.
The TGTSF task is defined by three principal inputs: the time series data, news messages, and channel descriptions. The challenge lies in the fact that news items are not directly linked to specific time series channels. Consequently, the model must infer the impact of each news item on individual channels based on their descriptions. This setup not only capitalizes on dynamic events but also integrates domain knowledge via channel descriptions. Such integration is crucial for the model to comprehend and learn the spatial correlations among channels, thereby enhancing the forecasting accuracy with nuanced contextual understanding.
Formally, the TGTSF task seeks to model the potentially complex joint distribution of future values in a multi-channel sequence. This modeling involves the integration of historical time series data, textual information from news messages, and channel descriptions over a specified look-back window:
$P(X_{n:n+h}^{j}|(X_{n-L:n-1}^{j},News_{n:n+h},Des^{j})),j\in[1,c]$ (1)
where $L$ denotes the look-back window length, $h$ stands for the prediction length, the $j$ means the channel number, $c$ is the total number of channels.
4 TGForecaster: A Baseline Model for The TGTSF Task
To validate the TGTSF task, we developed TGForecaster, a streamlined transformer model designed for multimodal fusion with cross-attention. Illustrated in Fig. 2, this model harnesses textual information to enhance the accuracy and relevance of time series forecasting. It employs a reimplemented PatchTST encoder for processing time series data and leverages off-the-shelf, pretrained text models for pre-embedding textual inputs into vector sequences across the time dimension, allowing for effective modality fusion in the embedding space.
Figure 2: Architecture of the TGForecaster Model. TGForecaster integrates three primary inputs: time series data from a look-back window, news embeddings, and channel description embeddings. The texts are aligned to the forecasting horizon length. Text-side embeddings act as queries within the modality mixer layer to synchronize outputs with the forecast length. On the right panel, TGForecaster’s efficiency is highlighted by the text encoder’s cross-attention layer, which effectively associates news with specific channels. The token-wise decoder prevents overfitting in the final linear layer through inverse patching. Key parameters—number of patches ($PN$), maximum news items per batch ($M$), look-back window ($L$), and forecast horizon ($H$)—are streamlined for clarity.
A key innovation in TGForecaster is the integration of a cross-attention layer to achieve the ’text guided channel independent’. This layer is pivotal in fusing news content with channel descriptions. Taking cues from the attention mechanism’s original use in recommendation systems, we treat the news as both the key and value, with the channel descriptions serving as queries. Through this mechanism, we calculate the relevance of each news item to every channel. The resulting scores are then used to aggregate the news, culminating in a composite embedding for each channel. Another crucial aspect of our design is the modality fusion, achieved through a separate layer of cross-attention where text embeddings function as query, while temporal embeddings act as the key and value. This ensures that the output dimensions align naturally with the query which has the length of forecasting horizon. The attention maps for these two cross-attention layers are provided in Appendix E. These maps compellingly demonstrate that the layers perform in accordance with our design specifications.
As decoder, we use the token-wise decoding to avoid the overfitting on the last layer as observed in previous work (Lee et al., 2023). Each token is project back to its original patch length and weighted sumed to get the final forecasting result. Notably, TGForecaster does not incorporate the RIN. Our experiments, detailed in Appendix D, demonstrate that the inclusion of RIN degrades performance when sufficient information is already available.
This design strategy not only enables the integration of textual data into the time series forecasting model but also facilitates a direct comparison of the impact of text on forecasting accuracy.
5 TGTSF Benchmark Datasets
We have designed four TGTSF benchmark datasets across three categories: synthetic, captioned existing, and real-world datasets. We plan to release these TGTSF datasets along with captions, captioning scripts, and pre-embeddings, adhering to all applicable intellectual property constraints.
5.1 Synthetic Toy Dataset
The Synthetic Toy Dataset evaluates models’ ability to utilize textual information for time series forecasting. It features diverse patterns, including segments of sinusoidal waves with variations in frequency, amplitude, and trend. Textual descriptions resembling news reports precede each change point, outlining upcoming alterations to prompt the TGTSF model to adjust its forecasts accordingly. This setup tests the model’s effectiveness in mitigating distribution shifts and harnessing textual cues for improved forecasting accuracy.
5.2 Electricity-Captioned Dataset
The Electricity-Captioned Dataset builds on a commonly-used electricity utility dataset (Zhou et al., 2021) that tracks appliance usage in an office building, with daily patterns largely influenced by whether it is a workday. This dataset has been enhanced with straightforward, publically-known textual information such as the type of day (e.g., day of the week, public holiday, workday), using the channel name as the descriptor. This enhancement serves a dual purpose: to illustrate whether minimal textual data can improve prediction accuracy and to facilitate a direct performance comparison with traditional time series models.
5.3 Weather-Captioned Dataset
The Weather-Captioned Dataset is designed to overcome the limitations of commonly used datasets in time series forecasting (TSF), especially the lack of predictability and periodicity in variables like rainfall, wind speed, and direction. By incorporating external climate information, this dataset demonstrates how integrating textual data can address the challenges of information insufficiency in traditional TSF, enabling enhanced control over forecasting outcomes.
Originally limited to a single year (Zhou et al., 2021), we have expanded the weather dataset to encompass a decade of detailed weather data from 2014 to 2023, sourced consistently from the same weather station. The Weather-Captioned dataset now includes weather forecasting reports from a publicly accessible service, detailing the climate conditions in Jena, Germany, where the weather station is approximately located. These reports, updated every six hours and daily, cover a range of parameters including weather conditions, temperature, humidity, wind speed, and direction, providing a comprehensive overview for our analysis. We directly use the channel name as channel description.
Table 1: Example caption of the Weather-Captioned dataset.
Topic Example
Month & Time of the Day It’s the early morning of a day in January.
Overall Weather The current weather is clear.
Weather Trend in next 6h The weather is expected to remain clear.
Temprature Trend in next 6h The temperature is showing a mild drop.
Wind Speed & Direction There is Light Breeze from NNW.
Atmosphere Pressure Level The atmospheric shows Average Pressure.
Humidity Level The air is very humid.
To transform this quantitative data into actionable insights for TSF, we employed GPT-4 to summarize each forecasting report into a set of seven thematic sentences, as demonstrated in Tab. 1. GPT-4 is prompted to avoid specific numerical details from the original reports to prevent information leakage and to simplify the model’s learning process. To further enrich the dataset, GPT-4 generated three distinct summary versions for each report, resulting in $21873^{7}=$ 2187 possible unique textual captions for each time step. The full details of the prompt used for generating these summaries are provided in the appendix.
We divided this extensive dataset into two subsets for detailed analysis: Weather-Captioned-Medium, covering data from 2014 to 2018, and Weather-Captioned-Large, which includes the entire dataset spanning ten years. This dataset encompasses over 525,600 weather data records across 21 different channels, with more than 85,000 unique sentences. This vast corpus provides millions of potential training samples for captioned time series forecasting. Further information is detailed in the appendix.
5.4 Steam Game Dataset
[…]

### v3

@@ 1 Introduction
Time series forecasting (TSF) has witnessed significant progress, yet recent studies indicate diminishing returns: deep learning models [1, 2, 3] or even pretrained time series foundation models [4, 5, 6] now deliver only marginal performance gains over simple linear baselines [7, 8, 9].
This performance plateau arises primarily because traditional TSF methods rely exclusively on historical data, inherently adopting a problematic "self-stimulation" assumption—forecasting models depend solely on past observations while ignoring external interventions. In reality, time series often originate from dynamic systems that evolve not just from their previous state but also through external interventions. With a control-theoretic framework, we demonstrate that this modeling gap imposes an insurmountable barrier on forecasting accuracy. Recent studies have made preliminary yet promising attempts to incorporate both textual [10, 11, 12] and exogenous variables [13, 14] context for forecasting, though lacking rigorous theoretical grounding. Our analytical framework explicitly demonstrates that incorporating intervention-related context can lower forecasting error bounds.
Motivated by this insight, we propose Intervention-Aware Time Series Forecasting (IATSF), a novel forecasting paradigm that incorporates external interventions into conditional predictions. Since interventions may take diverse, often qualitative forms, we focus on textual data due to its ubiquity and ability to encode nuanced, non-quantifiable signals. By modeling interventions explicitly, IATSF reframes forecasting from correlation-based inference to dynamic system modeling, providing a principled framework for integrating textual context as supplementary intervention signals. This approach not only aligns forecasting with real-world system dynamics but also offers practical advantages in interpretability and adaptability.
Despite these theoretical advances, practical adoption remains challenging due to the lack of datasets and models that are compatible with intervention-aware forecasting. Existing multimodal time series forecasting (TSF) approaches often rely on large language models (LLMs) and datasets [15, 10] optimized for prompting rather than structured intervention modeling. Consequently, these datasets often have: (1) short horizons limiting meaningful intervention evaluation; (2) overly simplistic or ambiguous textual descriptions causing information leakage or irrelevance; and (3) poor temporal synchronization between textual and numerical data. To address these limitations and operationalize our theoretical insights, we introduce the Temporal-Synced IATSF benchmark, explicitly designed with leak-free textual interventions synchronized to extended, realistic forecasting horizons.
To demonstrate the effectiveness of intervention-aware forecasting, we propose FIATS (Forecaster for Intervention-Aware Time Series), a lightweight, LLM-free baseline model. FIATS uses semantically aligned textual embeddings and introduces a novel Channel-Aware Adaptive Sensitivity Modeling (CASM) mechanism guided by control theory. Additionally, a Causal Alignment Decoder with Channel-Aware Parameter Sharing (CAPS) explicitly aligns textual interventions with forecasting channels. Extensive experiments across synthetic, physics-based, and market datasets show that FIATS consistently outperforms state-of-the-art methods, with ablation studies confirming that performance improvements stem from explicit intervention modeling rather than model complexity.
In summary, our key contributions are:
  • A control-theoretic analysis reveals intrinsic forecasting barriers caused by the "self-stimulation" assumption (sole reliance on history) and shows intervention-aware modeling reduces error bounds.
  • Building on this analysis, we introduce IATSF, a paradigm that models time series with external interventions, bridging the gap between traditional TSF and real-world dynamic systems.
  • We operationalize IATSF with the Temporal-Synced IATSF benchmark and a LLM-free FIATS model, whose performance gains are shown to stem from principled intervention modeling, not architectural complexity.
2 Background and Motivation: TSF from System Analysis Perspective
Time series data are typically measurements of real-world dynamic systems whose behaviors are continually shaped by external events. However, conventional datasets and forecasting methods often rely exclusively on historical measurements, neglecting these influential external factors. For instance, the widely-used ETT dataset [16] records power load and oil temperature from electric transformers—both significantly impacted by external events, including human activities and environmental conditions. Traditional approaches incorporating numeric exogenous variables, such as ARIMAX [17], have advanced forecasting capabilities by explicitly including external numeric inputs. Nevertheless, these methods fall short when dealing with qualitative, uncertain external factors frequently represented in textual form—such as event descriptions, news reports, or expert narratives. Recent works have attempted to bridge this gap by incorporating textual contextual information to improve forecasting accuracy [10, 11, 12], though a rigorous theoretical justification for their effectiveness remains lacking.
To systematically address this qualitative gap, we formally identify and analyze the intrinsic limitations of ignoring qualitative external interventions from a dynamical systems perspective [^1] All proofs and discussion are provided in Appendix B, unless otherwise specified..
2.1 Time Series Are Observation of Real-World Dynamic Systems
Consider a general dynamical system characterized by hidden states $Z\in\mathbb{R}^{m}$, evolving based on historical states and independent external interventions [18, 19, 20]:
$Z_{f}=F(Z_{h},U_{t}),\quad X=O(Z)$ (1)
where $F$ represents the true system dynamics, $U_{t}$ denotes time-varying independent external interventions, $O$ represents observation, $X$ for the the observed signal. For analytical clarity, we assume full observability, i.e. $X=Z$. We also discuss a simple linear system case $X_{f}=AX_{h}+BU_{t}$, where $A$ governs self-stimulated state transitions and $B$ encodes intervention sensitivity. Standard forecasting datasets $\mathcal{D}=\{(X_{h}^{(i)},X_{f}^{(i)})\}_{i=1}^{N}$ are generated through sliding window on the observed signals, where $X_{h},X_{f}$ stand for look-back window and forecasting horizon segment accordingly.
2.2 The Implicit Self-Stimulation Assumption in TSF
Traditional forecasting adopts a self-stimulation paradigm where models $f_{\theta}$ attempt to approximate system dynamics using only historical observations:
$f_{\theta}^{*}=\arg\min_{\theta}\mathbb{E}\left[\|\epsilon\|^{2}\right]=\arg\min_{\theta}\mathbb{E}\left[\|F(X_{h},U_{t})-f_{\theta}(X_{h})\|^{2}\right]$ (2)
The critical limitation stems from implicitly treating unobserved interventions as hidden random variables $U_{t}\sim\mathcal{P}_{U}$. This induces an irreducible forecasting error, as formalized by our first proposition:
Proposition 2.1 (Self-Stimulation Error Bound).
For any self-stimulated model $f_{\theta}$, it converges to predicting conditional expectation $F^{*}(X_{h},\mu)\triangleq\mathbb{E}_{U}[F(X_{h},U)]$, the prediction error covariance satisfies:
$\text{Cov}(\epsilon)\succeq\mathbb{E}_{X_{h}}\left[\nabla_{U}F\Sigma(\nabla_{U}F)^{\top}\right]$ (3)
where $\mu=\mathbb{E}(U_{t}),\quad\Sigma=\text{Cov}(U_{t})$. For linear systems, this falls back to:
$\text{Cov}(\epsilon)\succeq B\Sigma B^{\top}$ (4)
Proposition 2.1 reveals two fundamental limitations: 1) Self-stimulated models converge to predicting conditional expectations, rather than true dynamics, explaining prevalent averaging effects in practice as shown in Fig. 1, and 2) An irreducible error floor exists due to intervention stochasticity. This establishes a theoretical performance ceiling for conventional TSF approaches.
Figure 1: The real system runs under various interventions. The Intervention-Aware method can effectively approximate the real system according to the dataset while traditional self-stimulated method can only approximate a average scenario with persistent error, lead to bad or even collapse result. The right panel shows visualization result of a frequency modulated system which is very sensitive to the intervention, i.e. large $\nabla_{U}F$.
3 IATSF: Intervention-Aware Time Series Forecasting
3.1 Task Formulation
We propose Intervention-Aware Time Series Forecasting (IATSF) to overcome the self-stimulation limitation. The key innovation lies in explicit intervention modeling:
$f_{\theta}^{*}=\arg\min_{\theta}\mathbb{E}\left[\|F(X_{h},U_{t})-f_{\theta}(X_{h},U_{t})\|^{2}\right]$ (5)
where $U_{t}$ represents measurable interventions. This paradigm enables breaking the error bound in proposition 2.1 through intervention-aware learning, as detailed in Fig. 1.
As shown above, instead of assuming the external intervention stays the same in the TSF, IATSF aims to predict a conditioned future with the observed or predicted intervention even though it is not fully observed or precise. The error reduction mechanism is formalized through our second proposition:
Proposition 3.1 (Partial Intervention Efficacy).
For a system with $p$ independent interventions $U_{t}=\sum_{i=1}^{p}U_{t}^{i}$, incorporating any known intervention $U_{t}^{j}$ reduces the error covariance by:
$\Delta\text{Cov}(\epsilon)=\nabla_{U_{j}}F\Sigma_{j}(\nabla_{U_{j}}F)^{\top}$ (6)
For linear systems, this reduces the lower bound by $B_{j}\Sigma_{j}B_{j}^{\top}$.
Proposition 3.1 demonstrates that any measurable intervention information reduces forecasting uncertainty, even with incomplete intervention knowledge. This motivates our key insight: textual descriptions of interventions provide viable information for uncertainty reduction, despite non-numeric formats.
3.2 Language as an Intervention Modality
Incorporating exogenous variables is a common approach [13, 14], but it typically requires numerical time series or one-hot encoded inputs sampled at the same rate as the target series—even when the actual interventions are sparse. This limits flexibility, especially when new events occur. In real-world settings, many impactful factors—such as weather anomalies, geopolitical shifts, or human decisions—are hard to quantify but still essential for accurate forecasting. To address this, we propose modeling interventions using linguistic descriptors, which naturally capture compositional and relational semantics through lexical encoding. This allows for expressive representations of complex events (e.g., "simultaneous port strikes and agricultural subsidies") without incurring combinatorial overhead. This design offers several key advantages:
Expert Knowledge Integration: Textual interfaces facilitate the direct inclusion of domain-specific expertise via natural language specifications (e.g., "anticipated regulatory changes will suppress industrial output"). This makes it easier to incorporate human input or LLM-driven forecasting through linguistic conditioning of interventions.
Generalizability: Textual representations provide flexibility across various contexts, allowing models to generalize more effectively to new or unseen intervention scenarios. The use of natural language reduces reliance on rigid, pre-encoded numerical data, enabling better adaptability to diverse situations.
Cross-Modal Causal Alignment: By embedding both linguistic intervention descriptors and their temporal effects in a shared space, neural architectures can learn latent mappings that align interventions with their causal impacts on the system.
4 IATSF Benchmark Datasets
4.1 Leak-Free Dataset Design
The IATSF benchmark is explicitly constructed to be leak-free, adhering to the principle that models must not access future system states. To enforce this, we only include independently evolving interventions—external causal factors that influence the system but are not themselves outcomes of it. Including variables that directly describe or summarize the time series trajectory (as in [15, 3]) would violate this principle by introducing future state information; see Appendix O for further discussion.
Since system responses to interventions often occur much faster than the sampling interval (e.g., photovoltaic panels react to sunlight in milliseconds), we assume interventions take effect instantaneously and denote the up-to-date intervention as $U_{f}$. However, in real-world deployment, such ground-truth interventions are unavailable at prediction time. Therefore, we restrict the intervention input to three categories: (1) Known information, such as holidays or other common knowledge; (2) Predictions of $U_{f}$ derived from sources with expert knowledge, such as weather reports; and (3) Hypothetical or controlled events, which allow the IATSF models to simulate "what-if" scenarios during decision-making. Evaluation strategies accounting for prediction errors in interventions are detailed in Appendix B.3.
4.2 Brief Datasets Introduction
Each instance in IATSF is defined as $\mathcal{D}=\left\{((X_{h}^{(i)},U_{f}^{(i)},D),X_{f}^{(i)})\right\}_{i=1}^{N}$, comprising historical time series $X_{h}$, future-aligned interventions $U_{f}$, with $D$ denoting channel descriptors and future values $X_{f}$ as ground truth. $U_{f}$ may contain multiple temporally-aligned interventions, each independently observed alongside the time series. The datasets contains multiple channels or instances, each with distinct distributions.
As mentioned earlier, the key challenge is to identify time-synced interventions that are independently observed alongside the time series. This makes it impractical to apply traditional datasets, such as ETT [16], which lack the necessary contextual information. To address this, we have designed four initial IATSF benchmark datasets across four domains: synthetic controlled systems, physics, building management, and market analysis. Each dataset includes temporally aligned interventions and time series data, making them suitable for IATSF validation. Please check Appendix P for detail.
Frequency Modulated Toy Dataset
A simple intervention-aware system with sinusoidal wave segments and varying frequencies under the influence of interventions. Textual descriptions precede each change point, providing a clear context for upcoming alterations. With full intervention observation, this dataset has a theoretical error lower bound of 0.
Electricity Utility Dataset
Based on a widely used dataset for office building appliance usage [16], this dataset incorporates daily patterns affected by workdays. We enhance it with textual information (e.g., day type, public holidays), using channel names as descriptors. This dataset allows us to explore the impact of minimal textual data on prediction accuracy and compare it to traditional models.
Atmospheric Physics Dataset
Sourced from a research initiative monitoring fine-grained atmospheric signals. In addition to commonly reported variables such as temperature and humidity, it includes measurements like Dew Point and Short-Wave Downward Radiation (SWDR), which are closely linked to weather interventions but are excluded from general weather reports. This dataset observes an ideal system for studying IATSF, as it provides clean, direct, and highly correlated intervention effects. We use open-source weather report APIs to prepare textual interventions, thereby avoiding direct access to the time series data. By incorporating limited system-level predictions, this dataset demonstrates how expert knowledge and external information can improve the accuracy of fine-grained time series forecasting. It also includes multiple channels with distinct distributions and intervention responses, presenting challenges in channel-specific behavior modeling.
Game Active User Dataset (GAUD)
A key application of IATSF is to model the impact of business decision to the market. This dataset tracks daily active users for 90 games on an online platform, with developer, category, and update logs as interventions. It helps evaluate the model’s ability to capture market response to human interventions. The dataset is highly random with complex nonlinearities and includes short time series, allowing for evaluation in cold-start or zero-shot scenarios.
5 FIATS: A Simple System-Aware Baseline Model for IATSF
While recent studies [11, 10, 15, 12, 21] explore text-informed forecasting using the reasoning capabilities of large language models (LLMs), these approaches are limited to short sequences and simplistic interventions. They often suffer from high variance, low interpretability, and significant computational and token overhead. To address these limitations and rigorously validate the IATSF task, we introduce FIATS—the first LLM-free, numerical-based forecaster designed for intervention-aware time series. As illustrated in Fig. 2, FIATS combines a patch-based time series encoder [1] with novel text-embedding-based [22, 23] intervention semantic encoder and decoder. The novelty is as follows:
Temporal-Synced Intervention
Real-world systems often respond rapidly to interventions, necessitating temporal alignment between text and time series data. FIATS addresses this by synchronizing each time series patch with the last intervention observed, e.g. for patch start from 10:15, sync with the last timestep with intervention update of 10:00. This ensures the model uses only leak-free, contemporaneous interventions when forecasting subsequent patches, preventing future information leakage while maintaining temporal relevance.
Channel-aware Adaptive Sensitivity Modeling (CASM)
In FIATS, we reframe the attention mechanism through a control-theory perspective to explicitly model intervention sensitivity—a novel approach within attention-based architectures. Starting from linear systems where time series are observed by $X_{f}=CZ_{f}=CAZ_{h}+CBU_{f}$, channel-specific sensitivity to interventions is governed by $\frac{dx_{f}^{i}}{dU_{f}}=c^{i}B$. This indicates that each channel responds differently to external interventions. The error analysis is discussed in Appendix B.4. To capture this without introducing excessive parameters, we reconceptualize cross-attention as a Channel-aware Adaptive Sensitivity Modeling Block, as shown in the right panel of the Fig. 2, specifically:
- Query as Channel-wise Sensitivity $\tilde{C}=Desc\cdot W_{Q}$: Channel descriptions $Desc\in\mathbb{R}^{CN\times D}$ are served as query ($CN$ as channel number). The query projection explicitly learns how textual channel features (e.g., "atmospheric pressure") influence intervention sensitivity for each channel. This allows the model to adjust how interventions are perceived based on channel-specific characteristics.
- Key as Intervention Filter $\tilde{B}_{U_{f}}=(News\cdot W_{K})^{\top}$: The key projection maps temporal-synced news embeddings $News\in\mathbb{R}^{M\times D}$ to a system sensitivity matrix ($M$ as news number), allowing the model to filter out irrelevant interventions (e.g., excluding "tech stock news" when forecasting atmospheric physics). This ensures that only pertinent interventions are considered for each system.
- Value as Intervention Translator $\tilde{U_{f}}=News\cdot W_{V}$: Value projection learns to maps news text embedding to $\tilde{U_{f}}$, the latent space of actionable intervention effects.
Figure 2: Architecture of FIATS. FIATS integrates three inputs: time series data from a look-back window, temporal-synced news embeddings, and channel description embeddings. The intervention encoder employs CASM blocks in a residual connection along with multiple self-attention layers to enhance feature extraction. The CAPS causal alignment decoder projects the historical time series embeddings into the future, guided by channel-aware, time-synced interventions. A token-wise decoder is used to prevent overfitting in the final linear layer, as discussed in [24].
The above analysis show that the attention mechanism can effectively generate the channel-aware intervention $U_{f}^{c}$. This design allows identical interventions to differentially impact channels based on their descriptions. Unlike static sensitivity coefficients found in classical systems, this formulation maintains the nonlinear characteristics provided by the transformer block, allowing for greater learning flexibility to approximate complex nonlinear system. Additionally, it aligns well with the theoretical framework, making the model more interpretable. The attention map produced by the CASM layer directly reveals the sensitivity of each channel to various interventions, providing clear insights into how interventions impact different channels based on their specific descriptions.
Channel-Aware Parameter Sharing (CAPS)
While CASM addresses heterogeneous intervention responses, channels also exhibit inherent differences in their temporal patterns – a critical factor neglected by conventional parameter sharing. Previous shared models approximate all channels with a same set of parameters introducing persistent errors $\epsilon_{i}=o_{i}(Z)-\frac{1}{k}\sum_{j=1}^{k}o_{j}(Z)$ where $o_{i}$ for real system channel-specific dynamics.
To mitigate this issue, FIATS introduces a lightweight channel-aware decoding mechanism. All channels are first encoded into a shared latent space $\tilde{Z}$ by a unified time-series encoder. Then, a channel-conditioned decoder is used to adaptively project this latent representation into a channel-aware space, conditioned by the channel-specific time-synced intervention embeddings $U_{f}^{c}$. decoder approximates channel-specific adjustments by modulating the shared latent space through cross-attention $Attention(Q=U^{c}_{t},K,V=\tilde{Z})$ to simulate such nonlinear projection. To avoid future information leakage, we apply causal attention mask here. We will omit the analysis.
[…]